OpenStudy (anonymous):

Brian is creating a collage on a piece of cardboard that has an area of 120r^3 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of sqrt(r^5) square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (btaylor):

Divide the cardboard area by the area of each piece. 120r^3/sqrt(r^5). sqrt(r^5) = r^2.5, and r^3/r^2.5 = r^.5 so the answer is 120r^.5, or 120sqrt(r).

OpenStudy (anonymous):

Can you also do 20z/sqrt(z^13)

OpenStudy (btaylor):

ok...sqrt(z^13) = z^6.5. z/z^6.5 = z^-5.5. So the answer is 20(z^-5.5)